# Cosmic observation of a model in the horizon of $f(Q,C)$-gravity

### v1

@@ III.3 BAO Dataset
The exploration of baryonic acoustic oscillations occurs in the early universe, where baryons and photons are tightly coupled due to Thomson scattering. Because of the intense photon pressure, both baryons and photons behave as a unified fluid that resists gravitational collapse. Instead, they undergo oscillations, leading to the phenomenon known as Baryonic Acoustic Oscillations (BAO). BAO involves analyzing the spatial distribution of galaxies to measure how cosmic structures evolve during the expansion of the cosmos [82, 83, 84, 85, 86, 87]. This distinction provides a means to differentiate between various types of Dark Energy (DE). The sound horizon $r_{s}$ at the epoch of photon decoupling $z^{*}$, which defines the characteristic scale of BAO, is given by:
@@ IV.1 Cosmological Parameters
In the evolution of the Universe, we have discussed various physical and kinematic parameters whose characteristics can be explored either by analyzing their formulas or interpreting graphical representations. In this section, we have examined the formulae of some essential parameters like the deceleration parameter ($q$), jerk parameter ($j$), and EoS parameter ($\omega$) etc. The deceleration parameter can be computed by the formula $q=-1-\frac{\dot{H}}{H}$. Using the Eq. (20), $q$ can calculated as
@@
To study the evolution of the deceleration parameter, the graphical analysis is the most effective way and so in Fig. 6a, we plot the trajectories of $q$ using various latest observational datasets. In fig. 6a, we observe that the model transits from the early decelerating to the late-times accelerating state which is consistent according to the standard cosmology. Also, using the deceleration parameter we can find another kinematic parameter named jerk parameter $j$ given as
@@
In Fig. 6b, the evolution of the jerk parameter $j$ is observed. In this model, the trajectories of the jerk parameter are more consistent for the joint dataset OHD + Pantheon + SH0ES and for the other dataset, it deviates from the $\Lambda$CDM on high redshift $z$. However, the trajectories of the jerk parameter converge to the $\Lambda$CDM where $j=1$ as redshift $z\to-1$. In this model, we use the specific form of $f(Q,C)$ function given in the Sec. II. The energy density $\rho$ and the isotropic pressure $p$ are evaluated from the field equations (17-18) and the Eqs. (20-21). Since the expressions of the energy density and the isotropic pressure are quite large, therefore their evolutions have been studied by the depicted plots. In Fig. 7, the energy density and pressure both decrease monotonically from the evolution of the early Universe to later times. The value of $\rho$ is positive for the entire redshift range $z$, whereas $p$ becomes negative in the redshift range $z<1.5$, and can be seen in Figs. (7a-7b).
The energy density $\rho$, the pressure $p$, and the EoS parameter $\omega$ of the model play a crucial role in cosmology to understand the physical behavior of the universe. The equation of state is related by the equation $p=\omega\rho$. In literature, the equation of state (EoS) parameter tracks the evolution of the Universe through various stages, which begin from an early deceleration stage to the late-times acceleration stage. These stages encompass periods such as the stiff fluid epoch ($\omega=1$), radiation epoch ($\omega=\frac{1}{3}$), dust epoch ($\omega=0$), quintessence dark energy ($-1<\omega<-\frac{1}{3}$), the $\Lambda$CDM ($\omega=-1$), and the phantom dark energy ($\omega<-1$). The quintessence dark energy is stable whereas the phantom dark energy is highly unstable.
In this model, the behavior of EoS can be analyzed in Fig. 7c. Here, we observe the different stages of the Universe’s expansion to varying epochs in different ranges of the redshift. At present our model is in the quintessence region and in late times the value of $\omega$ converges to $-1$ which represents the $\Lambda$CDM model.
@@ IV.2 Energy Conditions
The Energy conditions (ECs) are the essential features of the model to understand the universe on a large scale. These energy conditions are essential for understanding the universe on a large scale. The energy limits are essential to explain the existence and formation of the matter. They help to characterize the ordinary and exotic matter distributions and are essential for confirming the validity of the theoretical framework.
These energy conditions have been deduced from the considerations of the Raychaudhuri equation which explains the dynamics of the expansion of the late-time evolution of the Universe. It is remarkable to note that these equations are feasible to discuss in physical, geometrical, and perfect fluid form according to Table 3 and do not explicitly reference gravitational field equations. The criterion for gravitational attraction simplifies to $R_{ij}k^{i}k^{j}\geq 0$. In general relativity, this criterion is expressed in terms of the stress-energy tensor $T_{ij}$ as $T_{ij}k^{i}k^{j}\geq 0$. In modified gravity theories, the field equations take the form $G_{ij}\equiv R_{ij}-\frac{1}{2}Rg_{ij}=T_{\text{eff},\mu\nu}$. Consequently, the condition $R_{ij}k^{i}k^{j}\geq 0$ leads to the null energy condition (NEC) in the modified context as $T_{\text{eff},ij}k^{i}k^{j}\geq 0$.
@@ IV.4 Perturbation Analysis
Where, $H_{0}(t)$, $\rho_{0}(t)$, $\delta_{H}(t)$, and $\delta_{\rho}(t)$ stand for the baseline Hubble parameter, baseline energy density, and perturbation parameters for the Hubble parameter and energy density, respectively. u Within the domain of matter fields, we confront equations of first-order perturbations. The particular equation employed in this model is
@@
where $b=K\rho_{m0}$ which is normalized to unit in this work. Here, the fluctuations in matter dictate the total perturbation around a cosmological solution (or vice versa) in the context of General Relativity (GR). The first-order perturbation equation employed in this model can now be reformulated after eliminating $\delta_{H}(t)$ using Eqs. (37) and (38) as
@@
To study the perturbation parameters, we have plotted the trajectories of $\delta_{\rho}(z)$ and $\delta_{H}(z)$ for all the observational datasets. In Fig. 10, we notice that the variation of perturbation parameters attain their small values throughout the whole range of $z$ and this indicates the stability in the model.
@@ V Conclusions
In the symmetric teleparallel approach to gravity, non-metricity takes precedence over curvature as the fundamental geometric quantity. Unlike the Levi-Civita Ricci scalar $R^{*}$, the non-metricity scalar $Q$ differs by an additional term $C$ that represents a total divergence term as $R^{*}=Q+C$. At the level of equations, general relativity, and symmetric teleparallel general relativity are equivalent. However, modifications involving functions of $R^{*}$ and $Q$ are not equivalent due to their distinct differential properties. In our study, we have studied $f(Q,C)$ gravity and cosmology by incorporating both $Q$ and the boundary term $C$ into the Lagrangian formulation. Initially, we derived the general field equations and subsequently applied them to the flat Friedmann-Robertson-Walker (FRW) metric within a cosmological context. By exploring connection choices, where our analysis uncovered an effective dark-energy component stemming from geometric considerations, which indicates to yield intriguing cosmological phenomena. Furthermore, our framework demonstrated the potential for an effective interaction between the matter and the dark energy.
@@
Using a specific form of the function $f(Q,C)$, we have illustrated how our model mirrors the conventional thermal history of the universe, encompassing epochs dominated by both matter and dark energy as required. Moreover, we have parameterized the scale factor and examined the behavior of dynamical and physical cosmic parameters by calculating the best-fit values of the model parameters. To obtain the optimal values of the model parameters $\alpha$, $\beta$, and the Hubble parameter $H_{0}$, we use the Markov Chain Monte Carlo (MCMC) method by applying the emcee codes in Python using various observational datasets OHD, Pantheon, SH0ES, BAO, and their joint datasets discussed in the Figs. 1-4. The posterior distribution of the reconstructed deceptive $f(Q,C)$ model under $1\sigma$ and $2\sigma$ CLs and the error bar plots of the Hubble parameter $H(z)$, distance modulus $\mu(z)$ exhibit the degree of consistency between our model and the $\Lambda$CDM in the standard cosmology (see Figs. 5).
@@
The current values of the physical parameters $H$ and $q$ and the model parameters for different observational datasets are accessed in Table 1. We study the behavior of the trajectories of the deceleration parameter $q$ for the best-fit values in Fig. 6a where the transition from the deceleration to acceleration occurs in the redshift range $20.70<z_{tr}<$ 2. At present, our model is in the accelerating state of the universe. The jerk parameter $j$ is positive for all the redshift ranges and converges to $1$ as $z\to-1$ for all the constrained model parameters shown in Fig. 6b. The evolution of the energy density is always positive and decreases as the cosmic redshift $z\to-1$ (see Fig. 7a).
@@
The isotropic pressure of the model is positive for the redshift range $z>1$ and shows the negative behavior for all the constraints of the model parameters as $z\to-1$. The model indicates the existence of dark energy in the universe that leads to the accelerating expansion of the universe in later times (see Fig. 7b). The EoS parameter transits from the perfect fluid state to the dark energy state which approaches $-1$ shown in Fig. 7c. Hence the the model shows the quintessence-like dark energy model and converges to the $\Lambda$CDM as $z\to-1$ in the standard cosmology. In fig. 8, It is observed that the model satisfies the energy conditions NEC and DEC but does not satisfy SEC in the later times indicating the existence of dark energy in late times.
@@
Fig. 9 depicts the behaviors of the Statefinder Diagnostic pairs $\{s,r\}$ and $\{q,r\}$. The model starts with the Chaplygin region, enters into the quintessence region, and approaches the $\Lambda$CDM in the parametric space $\{s,r\}$. The model transits from the deceleration zone to the acceleration zone, deviates from SCDM, crosses the Quintom line, and approaches the de Sitter state in the parametric space $\{q,r\}$. In Fig. 10, we have plotted the trajectories of $\delta_{\rho}(z)$ and $\delta_{H}(z)$ for all the observational datasets. It has been noticed that the variation of perturbation parameters attain their small values throughout the whole range of redshift $z$ which indicates the stability in the model.
@@
Finally, given the evolution of the various trajectories of the jerk parameter, the EoS parameter, and the statefinder of the model we can see that our model converges to the $\Lambda$CDM. Moreover, to examine the stability of the model we discuss the perturbations both in the Hubble parameter and the energy density. Hence, we conclude that the model is an accelerated expanding model with quintessence-like behavior at later times $z\to-1$.
@@ Acknowledgement
The authors J. K. Singh and Shaily thank the Dept. of Mathematics, NSUT, New Delhi-78, India, and Bennett University, Greater Noida, India for providing the necessary facilities where a part of this work has been completed. JRLS would like to thank CNPq (Grant no. 309494/2021-4), and PRONEX/CNPq/FAPESQ-PB (Grant nos. 165/2018, and 0015/2019) for financial support.

### v2

@@ III.3 BAO Dataset
The exploration of baryon acoustic oscillations occurs in the early Universe, where baryons and photons are tightly coupled due to Thomson scattering. Because of the intense photon pressure, both baryons and photons behave as a unified fluid that resists gravitational collapse. Instead, they undergo oscillations, leading to the phenomenon known as Baryon Acoustic Oscillations (BAO). BAO involves analyzing the spatial distribution of galaxies to measure how cosmic structures evolve during the expansion of the cosmos [82, 83, 84, 85, 86, 87]. This distinction provides a means to differentiate between various types of Dark Energy (DE) models. The sound horizon $r_{s}$ at the epoch of photon decoupling $z^{*}$, which defines the characteristic scale of BAO, is given by:
@@ IV.1 Cosmological Parameters
In the evolution of the Universe, we have discussed various physical and kinematic parameters whose characteristics can be explored either by analyzing their formulas or interpreting graphical representations. In this section, we have examined the formulae of some essential parameters, for instance, the deceleration parameter ($q$), jerk parameter ($j$), and EoS parameter ($\omega$). The deceleration parameter can be computed by the formula $q=-1-\frac{\dot{H}}{H}$. Using the Eq. (20), $q$ can is given by
@@
Graphical analysis is the most effective way to study the evolution of the deceleration parameter. In Fig. 6a, we plot the trajectories of $q$ using the latest observational datasets. In fig. 6a, we observe that the model transits from the early decelerating to the late-times accelerating state. Such a behavior is consistent with the standard cosmology. Also, from the deceleration parameter, we can find another kinematic parameter named jerk parameter $j$, whose equation is
@@
In Fig. 6b, the evolution of the jerk parameter $j$ is observed. In this model, the trajectories of the jerk parameter are more consistent for the joint dataset OHD + Pantheon + SH0ES, and for the other dataset, it deviates from the $\Lambda$CDM on high redshift $z$. However, the trajectories of the jerk parameter converge to the $\Lambda$CDM where $j=1$ as redshift $z\to-1$. In this model, we use the specific form of $f(Q,C)$ function given in the Sec. II. The energy density $\rho$ and the isotropic pressure $p$ are evaluated from the field equations (17-18) and the Eqs. (20-21). Since the expressions of the energy density and the isotropic pressure are quite large, their evolutions have been studied using the depicted plots. In Fig. 7, the energy density and pressure decrease monotonically from the evolution of the early Universe to later times. The value of $\rho$ is positive for the entire redshift range $z$, whereas $p$ becomes negative in the redshift range $z<1.5$, and can be seen in Figs. (7a-7b).
The energy density $\rho$, the pressure $p$, and the EoS parameter $\omega$ of the model play a crucial role in cosmology to understand the physical behavior of the Universe. The equation of state is given by the ratio between pressure and density, whose formula is $\omega=p/\rho$. In literature, the equation of state (EoS) parameter tracks the evolution of the Universe through various stages, which begin from an early deceleration stage to the late-times acceleration stage. These stages encompass periods such as the stiff fluid epoch ($\omega=1$), radiation epoch ($\omega=\frac{1}{3}$), dust epoch ($\omega=0$), quintessence dark energy ($-1<\omega<-\frac{1}{3}$), the $\Lambda$CDM ($\omega=-1$), and the phantom dark energy ($\omega<-1$). The quintessence dark energy represents a stable evolution of the Universe whereas the phantom dark energy is highly unstable.
In this model, the behavior of EoS can be analyzed in Fig. 7c, where we can see different eras that our Universe passes through, depending on the redshift. At the present time, our model is in the quintessence region, and in late times, the value of $\omega$ converges to $-1$, converging to the same behavior of the $\Lambda$CDM model.
@@ IV.2 Energy Conditions
The Energy conditions (ECs) are the essential features of the model to understand the Universe on a large scale, once the energy limits can explain the existence and formation of the matter. They help to characterize the ordinary and exotic matter distributions and can confirm the validity of the theoretical framework.
These energy conditions have been deduced from the considerations of the Raychaudhuri equation which explains the dynamics of the expansion of the late-time evolution of the Universe. It is remarkable to note that these equations are feasible to discuss in physical, geometrical, and perfect fluid form according to Table 3 and do not explicitly reference gravitational field equations. The criterion for gravitational attraction simplifies to $R_{ij}k^{i}k^{j}\geq 0$. In general relativity, this criterion is expressed in terms of the stress-energy tensor $T_{ij}$ as $T_{ij}k^{i}k^{j}\geq 0$. In modified gravity theories, the field equations take the form $G_{ij}\equiv R_{ij}-\frac{1}{2}Rg_{ij}=T_{\text{eff},\mu\nu}$. Consequently, the condition $R_{ij}k^{i}k^{j}\geq 0$ leads to the null energy condition (NEC) in the modified context as $T_{\text{eff},ij}k^{i}k^{j}\geq 0$ [90].
@@ IV.4 Perturbation Analysis
where, $H_{0}(t)$, $\rho_{0}(t)$, $\delta_{H}(t)$, and $\delta_{\rho}(t)$ stand for the baseline Hubble parameter, baseline energy density, and perturbation parameters for the Hubble parameter and energy density, respectively.
Within the domain of matter fields, we confront equations of first-order perturbations. The particular equation employed in this model is
@@
where $b=K\rho_{m0}$ which is normalized to unit in this work. Here, the fluctuations in matter dictate the total perturbation around a cosmological solution (or vice versa) in the context of General Relativity (GR). The first-order perturbation equation employed in this model can now be rewritten by taking $\delta_{H}(t)$ in eq. (38) and replacing it into (37), yielding to
@@
To study the perturbation parameters, we have plotted the trajectories of $\delta_{\rho}(z)$ and $\delta_{H}(z)$ for all the observational datasets. In Fig. 10, we notice that the variation of perturbation parameters attain their small values throughout the whole range of $z$ and this indicates the stability of the model.
@@ V Conclusions
In the symmetric teleparallel approach to gravity, non-metricity takes precedence over curvature as the fundamental geometric quantity. Unlike the Levi-Civita Ricci scalar $R^{*}$, the non-metricity scalar $Q$ differs by an additional term $C$ that represents a total divergence term as $R^{*}=Q+C$. At the level of equations, general relativity and symmetric teleparallel general relativity are equivalent. However, modifications involving functions of $R^{*}$ and $Q$ are not equivalent due to their distinct differential properties. In our study, we have studied $f(Q,C)$ gravity and cosmology by incorporating both $Q$ and the boundary term $C$ into the Lagrangian formulation. Initially, we derived the general field equations and subsequently applied them to the flat Friedmann-Robertson-Walker (FRW) metric within a cosmological context. By exploring connection choices, our analysis uncovered an effective dark-energy component stemming from geometric considerations, which yields intriguing cosmological phenomena. Furthermore, our framework demonstrated the potential for an effective interaction between the matter and the dark energy.
@@
Using a specific form of the function $f(Q,C)$, we have illustrated how our model mirrors the conventional thermal history of the Universe, encompassing epochs dominated by both matter and dark energy as required. Moreover, we have parameterized the scale factor and examined the behavior of dynamical and physical cosmic parameters by calculating the best-fit values of the model parameters. To obtain the optimal values of the model parameters $\alpha$, $\beta$, and the Hubble parameter $H_{0}$, we use the Markov Chain Monte Carlo (MCMC) method by applying the emcee codes in Python using various observational datasets such as OHD, Pantheon, SH0ES, BAO. The constraints and features are shown in detail in Figs. 1-4. The posterior distribution of the reconstructed deceptive $f(Q,C)$ model under $1\sigma$ and $2\sigma$ CLs and the error bar plots of the Hubble parameter $H(z)$, distance modulus $\mu(z)$ exhibit the degree of consistency between our model and the $\Lambda$CDM (see Figs. 5).
@@
The current values of the physical parameters $H$ and $q$ and the model parameters for different observational datasets are accessed in Table 1. We investigated the behavior of the trajectories of the deceleration parameter $q$ for the best-fit values in Fig. 6a where the transition from the deceleration to acceleration occurs in the redshift range $20.70<z_{tr}<$ 2. The deceleration parameter also corroborates with an accelerating expansion of our Universe at the present time. The jerk parameter $j$ is positive for all the redshift ranges and converges to $1$ as $z\to-1$ for all the constrained model parameters shown in Fig. 6b. The evolution of the energy density is always positive and decreases as the cosmic redshift $z\to-1$ (see Fig. 7a).
@@
The isotropic pressure of the model is positive for the redshift range $z>1$ and shows the negative behavior for all the constraints of the model parameters as $z\to-1$. The model indicates the existence of dark energy in the Universe that leads to the accelerating expansion of the Universe in later times (see Fig. 7b). The EoS parameter transits from the perfect fluid state to the dark energy state, approaching $-1$, as shown in Fig. 7c. Hence, the model unveils the quintessence-like dark energy model and corroborates with $\Lambda$CDM as $z\to-1$. In fig. 8, we can verify that the model satisfies the energy conditions NEC and DEC but does not satisfy SEC for later times, indicating the existence of a dark energy regime.
@@
Fig. 9 depicts the behaviors of the statefinder diagnostic pairs $\{s,r\}$ and $\{q,r\}$. By analyzing the parametric space $\{s,r\}$, we observe that the model starts in a Chaplygin region, enters the quintessence region, and then corroborates with $\Lambda$CDM. Therefore, the model transits from the deceleration to the acceleration regime, deviates from SCDM, crosses the Quintom line, and approaches the de-Sitter state in the parametric space $\{q,r\}$. In Fig. 10, we have depicted the trajectories of $\delta_{\rho}(z)$ and $\delta_{H}(z)$ for all the observational datasets. It has been noticed that the variation of perturbation parameters attain their small values throughout the whole range of redshift $z$, which indicates the stability in the model.
@@
Finally, given the evolution of the various trajectories of the jerk parameter, the EoS parameter, and the statefinder, we can see that our model corroborates with the $\Lambda$CDM. Moreover, in analyzing the stability of the model, we discuss the perturbations for both the Hubble parameter and the energy density. Hence, we conclude that the model reflects an accelerated expanding model with quintessence-like behavior at later times $z\to-1$.
It is relevant to point out that the techniques presented in this work can be applied to other models of gravity, such as in $f(R,T)-\Lambda(\phi)$ [108], $f(T,{\cal T})$ [109], $f(Q,T)$ [111], and Weyl $f(Q,T)$ [110]. Moreover, we can also investigate in detail the mapping of further data in low redshift ranges such as those from DESI [112] and also from new and future surveys like Euclid [113] and BINGO [114, 115] telescopes.
@@ Acknowledgement
The authors J. K. Singh and Shaily thank the Dept. of Mathematics, NSUT, New Delhi-78, India, and Bennett University, Greater Noida, India for providing the necessary facilities where a part of this work has been completed. JRLS would like to thank CNPq (Grant no. 309494/2021-4), and FAPESQ-PB (Grant 11356/2024) for financial support.